OpenStudy (anonymous):

36 ft 3 in. to 37 ft 6 in. Find the percent of change round to the nearest tenth if necessary

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@pgpilot326

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

halp

OpenStudy (anonymous):

((37.5-36.25)/36.25) 100 = 3.4%

OpenStudy (skullpatrol):

$$\frac{\text{percent of change}}{100} = \frac{\text{change in amount}}{\text{original amount}}$$